Random samples of size n=2 are drawn from a finite population consisting of the numbers 5,67,8 and 9. Compute the mean, variance and standard deviation of the population, and the me a variance and standard deviation of the sample means. Showtablr with complete solutions


1
Expert's answer
2021-05-25T09:01:52-0400

We have given the population,

"5,6,7,8 \hspace{2mm}and \hspace{2mm}9"

Population mean = "\dfrac{5+6+7+8+9}{5} = 7"


Variance "= \dfrac{(5-7)^2+(6-7)^2+(7-7)^2+(8-7)^2+(9-7)^2}{5}\\"


"= \dfrac{4+1+0+1+4}{5} = 2"


Standard Deviation "= \sqrt{2} = 1.41"

Table for sampling distribution can be formed as




Variance of sampling distribution =

"\dfrac{(5.5-7)^2+(6-7)^2+2(6.5-7)^2+2(7-7)^2+2(7.5-7)^2+(8-7)^2+(8.5-7)^2}{10}\\
 = \dfrac{2.25+1+0.5+0+0.5+1+2.25}{10}
\\ = 0.75"


Sample standard deviation "= \sqrt{0.75} = 0.86"